OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. y - 2 = 3(x - 5) y + 8 = 3(x + 5) y - 2 = 3(x - 1) y - 5 = 3(x - 2)

OpenStudy (anonymous):

I need the answer.

OpenStudy (anonymous):

EYYYYYYYYYYYYYYYYYYYY

OpenStudy (anonymous):

EYYYYYYYYYYYYYYYYYYYYYYYY HELP EYYYY CONTACT EYYYYYYYYYYYY!

OpenStudy (anonymous):

come on guys where are you all.

OpenStudy (whpalmer4):

Okay, point-slope form is: \[y-y_0 = m(x-x_0)\]for a line with slope \(m\) passing through point \((x_0,y_0)\) Simple fill-in-the-blanks work...

OpenStudy (anonymous):

I need the answer please.

OpenStudy (whpalmer4):

I told you what to do. If you need the answer, do it. I'll check your work and tell you if you got it correct.

OpenStudy (whpalmer4):

If you don't understand what to do, say so, and I'll try to help you understand.

OpenStudy (anonymous):

y - 2 = 3(x - 5) got it gg.

OpenStudy (whpalmer4):

No, I don't think that is correct. Remember, points are (x,y), not (y,x)... If your equation is correct, the following will be true: \[5-2=3(2-5)\]\[3=3(-3)\]\[3=-9\]